# Universal Few-Shot Spatial Control for Diffusion Models

### v1

@@ 1 Introduction
Text-to-Image (T2I) diffusion models Esser et al. (2024); Rombach et al. (2022); Podell et al. (2024); Saharia et al. (2022); Chen et al. (2024b, a); Labs (2024), trained on large-scale datasets, have achieved remarkable success in generating high-quality, semantically aligned images from natural language prompts. While language-based control offers intuitive and flexible guidance, it often lacks the precision needed for fine-grained visual control, such as specific object positions, shapes, or scene layouts. To overcome this, recent works Huang et al. (2023); Mou et al. (2024); Li et al. (2023b); Zhang et al. (2023); Li et al. (2024); Qin et al. (2023); Zhao et al. (2023); Wang et al. (2023) incorporate explicit spatial signals—like edge maps, depth maps, and segmentation masks to control diffusion models.
To enable spatial control while preserving the generative quality of pre-trained diffusion models, existing methods typically employ control adapters Zhang et al. (2023); Mou et al. (2024); Li et al. (2023b) that inject spatial signals into a frozen T2I model. However, these adapters are usually trained independently for each spatial control task, requiring substantial computational resources and extensive labeled data for a new task. Alternatively, reusing pre-trained multi-task adapters - either directly Qin et al. (2023); Wang et al. (2023) or with minimal updates Zhao et al. (2023)- struggle to generalize to tasks that differ from their training distribution, and often show poor adaptability.
@@
To overcome these challenges, UFC introduces a universal control adapter that represents novel spatial conditions by adapting the interpolation of visual features of images in a small support set, rather than directly encoding task-specific conditions. The interpolation is guided by patch-wise similarity scores between the query and support conditions, modeled by a matching module Kim et al. (2023). Since image features are inherently task-agnostic, this interpolation-based approach naturally provides a unified representation, enabling effective adaptation across diverse spatial tasks. Furthermore, to facilitate rapid and data-efficient updates, UFC combines episodic meta-learning on a multi-task dataset with a small set of task-specific parameters, enabling effective adaptation under few-shot settings.
Extensive experiments across diverse spatial control modalities, including edge maps, depth maps, normal maps, human pose, and segmentation masks for human body, demonstrate that UFC outperforms prior approaches in few-shot settings by large margins. Notably, UFC achieves fine-grained spatial control using only 30 shots, and even on par with some fully supervised baselines on Normal, Depth, and Canny with just 150 support examples (0.1% of the training data). Furthermore, UFC is compatible with both UNet Ronneberger et al. (2015) and DiT Peebles and Xie (2023) backbones, ensuring applicability across recent T2I diffusion architectures. To our knowledge, we propose the first method for few-shot spatial control in text-to-image diffusion models, which enables spatial control image generation with novel conditions at test time using minimal annotated data.
@@ 2 Related Works
Text-to-Image diffusion models Esser et al. (2024); Rombach et al. (2022); Podell et al. (2024); Saharia et al. (2022); Chen et al. (2024b, a); Labs (2024) have gained widespread adoption for their ability to generate high-quality, semantically aligned images from text prompts. To enable finer control over image structure, various methods Mou et al. (2024); Zhang et al. (2023); Li et al. (2023b, 2024); Qin et al. (2023); Zhao et al. (2023); Wang et al. (2023); Tan et al. (2024); Wang et al. (2025) proposed incorporating additional spatial control inputs such as edge maps, segmentation masks, or human poses, etc., into diffusion models. ControlNet Zhang et al. (2023) augments a pretrained T2I diffusion model with a control adapter, initialized from the model itself, to encode spatial conditions and inject them into the frozen diffusion model. Uni-ControlNet Zhao et al. (2023) extends this by adding lightweight feature extractors to support multiple condition types. Prompt Diffusion Wang et al. (2023), built on ControlNet, introduces a vision-language prompting mechanism to enable in-context learning for handling a diverse set of tasks. However, these prior methods are limited to tasks seen during training. To address this limitation, training-free methods Bansal et al. (2023); Yu et al. (2023); Mo et al. (2024); Lin et al. (2024) have been proposed; but, they often rely on a long process of latent optimization, which substantially increases generation time while providing only limited controllability. Consequently, few-shot adaptation with more fine-grain controllability without incurring significant generation-time overhead remains unexplored.
@@
To extend diffusion models to unseen tasks, several works Cao and Gong (2024); Zhu et al. (2022); Yang et al. (2024); Sinha et al. (2021); Giannone et al. (2022); Li et al. (2023a); Jin et al. (2025) have explored few-shot learning for image generation. For example, D2C Sinha et al. (2021) trains a variational autoencoder Kingma and Welling (2014) on a few labeled examples to produce latent representations that condition a diffusion model to generate images in the target domain. Similarly, FSDM Giannone et al. (2022) uses visual features extracted from a small support set to guide image generation toward unseen classes. While these approaches demonstrate promising few-shot generation capabilities, they primarily focus on domain or class transfer, and are not designed to generalize across diverse spatial control tasks.
@@
Several recent works Gu et al. (2024); Yang et al. (2023) leverage the analogy between a query source and a support source-target pair to enable generalization in diffusion models on diverse tasks with visual instruction. Analogist Gu et al. (2024) treats visual instruction as an inpainting task, arranging query-support pairs in a $2\times 2$ grid and using a diffusion model to complete the missing cell. This training-free method generalizes well to tasks like colorization, deblurring, and editing. While we also leverage query-support analogy in few-shot settings, our goal differs: prior work preserves the support image’s appearance, whereas we aim to generate diverse content guided by spatial conditions.
@@ 3 Background
A text-to-image diffusion model generates an image $x$ conditioned on a text prompt $c_{\text{text}}$ by iteratively denoising a sequence of latent variables. Specifically, the model learns a finite-length Markov chain $\{z_{T-t}\}_{t=0}^{T}$ that starts from standard Gaussian noise $z_{T}\sim\mathcal{N}(0,I)$, where each transition $z_{t+1}\to z_{t}$ progressively removes noise. The final clean variable $z_{0}$ can be either the image itself or a latent representation $z=E(x)$, which is decoded back to the image by $x=D(z)$, where $(E,D)$ denotes an auto-encoder Rombach et al. (2022); Podell et al. (2024). The chain is the time-reversal of a diffusion process $\{z_{t}\}_{t=0}^{T}$ that gradually corrupts the data by adding small Gaussian noise until it becomes pure noise at $z_{T}$ Sohl-Dickstein et al. (2015). To learn each denoising transition, the diffusion network $\mathcal{E}_{\phi}$ is trained to predict the added noise $\epsilon$ at timestep $t$:
@@
where ($z_{0},c_{\text{text}}$) is drawn from the training dataset $\mathcal{D}_{\text{train}}$ and $t$ is uniformly sampled from $\{1,\cdots,T\}$. $z_{t}$ can be computed in closed form from $z_{0}$ and $t$ using the properties of Gaussian variables Ho et al. (2020). After training, an image is generated by running the learned Markov chain from $z_{T}$ down to $z_{0}$.
Although text-to-image diffusion models control the overall content of an image, a text prompt alone is limited in steering the fine-grained spatial structures. Spatial control approaches Mou et al. (2024); Zhang et al. (2023); Li et al. (2023b); Tan et al. (2024) address this limitation by introducing an additional spatial condition $y_{\tau}$, typically a dense map describing the target image (e.g., edges, depth, or semantic masks). For each condition type (or task) $\tau$, an auxiliary control adapter $\mathcal{G}_{\theta_{\tau}}(y_{\tau})$ [^1] The control adapter can take auxiliary inputs such as noisy latent, timestep, and text prompt, but we omit them for clarity. is trained and injected into a frozen, pre-trained diffusion backbone $\mathcal{E}_{\phi}$, while keeping the backbone parameters $\phi$ fixed to preserve their learned priors:
@@
However, learning a dedicated adapter for each task $\tau$ usually requires thousands of labeled data to achieve reasonable performance Zhang et al. (2023). This can be costly if we want to condition the pre-trained model on a novel task $\tau_{\text{novel}}$. On the other hand, reusing an adapter trained on a fixed task(s) for novel tasks often suffers from limited adaptation performance, especially when the new tasks are substantially different from training Zhao et al. (2023). Therefore, designing a versatile control adapter that can accommodate arbitrary unseen spatial conditions with a small amount of data remains a key challenge for data-efficient steering of text-to-image diffusion models.
@@ 4.1 Problem Setup
Our goal is to build a universal few-shot framework that steers a frozen text-to-image diffusion model with arbitrary novel spatial conditions, given only a handful of annotated examples (typically a few dozen). To this end, we introduce a universal control adapter $\mathcal{I}(y_{\tau};\mathcal{S}_{\tau})$ that can inject any type of spatial conditions into the frozen diffusion model by directly incorporating the given data as follows:
@@
where $\mathcal{S}_{\tau}=\{(x^{i},y_{\tau}^{i})\}_{i\leq N}$ denotes a support set comprising $N$ image-condition pairs for task $\tau$. We focus exclusively on the design of the adapter $\mathcal{I}$ while the remaining components, such as loss function, diffusion architecture, and injection strategy, are indepedent to the adapter and can be freely chosen from existing spatial control approaches Mou et al. (2024); Zhang et al. (2023); Li et al. (2023b).
@@
    Heterogeneous input formats. Spatial conditions vary widely—from extremely sparse cues such as human-pose keypoints to dense maps such as depth or segmentation. The adapter must transform these disparate inputs into the unified control features that the frozen diffusion model can consume reliably.
@@
    Severe data scarcity. For a new task, we assume only a few dozen labeled examples. To handle distribution shifts without over-fitting the tiny support set, the adapter needs an adaptation mechanism that is both flexible and efficient.
@@ 4.2 Universal Few-Shot Control Adapter
The patch-wise matching formulation in Eq. 4, inspired by few-shot dense prediction methods Kim et al. (2023, 2024), offers several advantages for the universal control problem. First, since the task-specific conditions are only used to determine the weights for the task-agnostic visual features, the resulting control features $\mathcal{I}(y_{\tau}^{q};\mathcal{S}_{\tau})$ reside in a unified visual feature space regardless of the task $\tau$. This ensures consistent guidance of the diffusion model and robust generalization to unseen tasks. Second, by composing the control features patch-wise, it can exploit the locality inductive bias on the image-condition pairs and amplify the effective number of support patches (i.e., bases) for each query patch. This provides sufficient representational power to model complex spatial controls from just a few support images. Finally, the general matching rule can wrap any control adapter $G_{\theta_{\tau}}$ by treating it as a condition encoder $g_{\tau}$ and adding an image encoder $f$. UFC can therefore inherit advances in adapter architectures and injection mechanisms developed for spatial control.
To adapt rapidly to a new task without over-fitting, UFC introduces only a small set of task-specific parameters $\theta_{\tau}$ while sharing the remainder. Starting from a pre-trained visual backbone $g(\cdot;\theta)$, we obtain a condition encoder
@@
where $\theta$ is shared across tasks and $\theta_{\tau}$ is lightweight (e.g., bias Ben Zaken et al. (2022) or LoRA Hu et al. (2022) parameters). After meta-training on a fixed task set (Section 4.4), $\theta$ is frozen and only $\theta_{\tau}$ is fine-tuned for each novel task $\tau$. The image encoder $f$, also initialized from a pre-trained visual backbone, remains frozen throughout to preserve task-agnostic visual features. Leveraging strong priors of large-scale visual pre-training and modern parameter-efficient tuning techniques, UFC generalizes to unseen tasks with minimal task-specific parameters.
@@ 4.3 Architecture
The formulation of UFC’s control adapter in Eq. (4) is general and can be incorporated into various diffusion backbones and adapter architectures. In this section, we describe the instantiation of UFC within the ControlNet framework Zhang et al. (2023) due to its popularity and strong performance. For clarity, we consider the pre-trained Stable Diffusion Rombach et al. (2022); Esser et al. (2024) with UNet backbone Rombach et al. (2022) in this section, while the one with DiT backbone Esser et al. (2024) is described in Appendix D.
We initialize the condition encoder $g_{\tau}$ and image encoder $f$ as separate copies of the UNet encoder of the pre-trained diffusion model, where the image encoder is kept frozen. For the task-specific parameters $\theta_{\tau}$ of the condition encoder, we adopt bias-tuning Ben Zaken et al. (2022), which has been proven to be flexible for few-shot learning in dense visual prediction Kim et al. (2023, 2024). The matching module is implemented as a multi-head cross-attention block Vaswani et al. (2017a), where the timestep embedding is fused through adaptive layer normalization Perez et al. (2018); Peebles and Xie (2023) inside the attention mechanism. More implementation details about the matching module can be found in Appendix A.1.
@@ 4.4 Training and Inference
To equip UFC with few-shot learning capabilities, we adopt the standard episodic meta-training protocol, wherein each training iteration is an episode that mirrors the test-time procedure. In each episode, we first sample a specific control task $\tau$ from a meta-training dataset $\mathcal{D}{{}_{\text{train}}}$ that consists of multiple tasks. The sampled task $\tau$ supplies a support set $\mathcal{S}_{\tau}$ and a query set $Q_{\tau}$, where the model exploits the support set to control the diffusion model with the query condition. The model is optimized using the standard denoising loss Ho et al. (2020) (or its flow-matching variant Lipman et al. (2023)):
@@ 5.1 Experimental setup
For experiments using the UNet architecture, we adopt Stable Diffusion v1.5 Rombach et al. (2022) as the diffusion backbone. Spatial conditions and support images are first projected into the latent space using the pre-trained VAE Kingma and Welling (2014) from the diffusion model, then processed by respective encoders. When encoding these inputs, we fix the denoising timestep to zero, provide empty text prompts, and omit the noisy latent input in order to focus solely on spatial information. Detailed model setting using DiT backbone is provided in the Appendix D.
@@
To enable episodic meta-training, we sample 300K text-image pairs from LAION-400M Schuhmann et al. (2021), with 150K containing humans (for Pose and Densepose) and 150K randomly sampled. Images are resized to the resolution $512\times 512$. Spatial conditions are extracted using pretrained off-the-shelf models. We consider six condition modalities that cover diverse semantics: Canny edge Canny (1986), HED edge Xie and Tu (2015), MiDaS Depth and surface Normal Ranftl et al. (2020), human Pose Cao et al. (2017), and human Densepose Güler et al. (2018). We consider Densepose as a control signal for semantic segmentation to constrain all segmentation classes to be included in the support set. During the training process, we maintain the setting that the models will be trained on 150K image-condition pairs for each task.
@@
We train UFC (UNet diffusion backbone) for 12.5K iterations with a batch size of 96 on 8 NVIDIA RTX 3090 GPUs, using AdamW Loshchilov and Hutter (2019) with a learning rate of $1\times 10^{-5}$. During inference with the UNet backbone, we adopt the PNDM sampler Liu et al. (2022) with 50 denoising steps, classifier-free guidance (CFG) Ho and Salimans (2021) scale of 7.5, and seed 42. For testing, we fine-tune UFC using all image–condition pairs in the support set, but use only a subset of them as conditions during inference to accommodate memory constraints and reduce generation-time overhead.
@@
  • Fully-supervised baselines ControlNet Zhang et al. (2023) and Uni-ControlNet Zhao et al. (2023) are both trained in a fully-supervised manner on the entire dataset (150K images for each task). ControlNet is trained separately for each condition type, while Uni-ControlNet is jointly trained with all conditions.
  • Few-shot baselines As the first to tackle spatial control in diffusion models under a few-shot setting, we adapt existing methods, Uni-ControlNet and Prompt Diffusion Wang et al. (2023), to align with our few-shot adaptation. For Uni-ControlNet, as mentioned in the original paper, we extend the input layer with additional channels for novel conditions and few-shot fine-tune this layer. For Prompt Diffusion, we evaluate both the zero-shot setup as proposed in the original paper and the full fine-tuning setup. For fair comparison, these few-shot methods are pre-trained using the same meta-training dataset as our method and fine-tuned on the same few-shot data.
  • Training-free baselines We also compare our method with training-free methods, Ctrl-X Lin et al. (2024) and FreeControl Mo et al. (2024), which directly employ the frozen T2I diffusion models for spatial control. We consider Ctrl-X as our main evaluation but leave comparisons with FreeControl in Appendix B using a smaller evaluation set due to its slow inference speed.
@@
We quantitatively evaluate all methods on COCO 2017 Lin et al. (2014) validation split, which contains 5,000 images with associated captions. All images and their corresponding conditions are resized to $512\times 512$. For tasks Pose and Densepose, we restrict evaluation to images containing humans. To assess image generation quality, we report the Fréchet Inception Distance (FID) Heusel et al. (2017). To evaluate controllability, we extract spatial conditions from the generated images using the same pre-trained networks used for dataset construction and compare them with the query conditions using task-specific metrics. Specifically, we use the Structural Similarity Index (SSIM) for Canny and HED, Average Precision ($\text{AP}^{50}$) computed using Object Keypoint Similarity for human Pose, Mean Squared Error (MSE) for Depth prediction, Mean Angular Error (MAE) for surface Normal, and mean Intersection over Union (mIoU) for human segmentation in Densepose.

### v2

@@ 1 Introduction
Text-to-Image (T2I) diffusion models [11, 41, 38, 43, 8, 7, 25], trained on large-scale datasets, have achieved remarkable success in generating high-quality, semantically aligned images from natural language prompts. While language-based control offers intuitive and flexible guidance, it often lacks the precision needed for fine-grained visual control, such as specific object positions, shapes, or scene layouts. To overcome this, recent works [19, 35, 28, 58, 27, 39, 59, 53] incorporate explicit spatial signals—like edge maps, depth maps, and segmentation masks to control diffusion models.
To enable spatial control while preserving the generative quality of pre-trained diffusion models, existing methods typically employ control adapters [58, 35, 28] that inject spatial signals into a frozen T2I model. However, these adapters are usually trained independently for each spatial control task, requiring substantial computational resources and extensive labeled data for a new task. Alternatively, reusing pre-trained multi-task adapters - either directly [39, 53] or with minimal updates [59]- struggle to generalize to tasks that differ from their training distribution, and often show poor adaptability.
@@
To overcome these challenges, UFC introduces a universal control adapter that represents novel spatial conditions by adapting the interpolation of visual features of images in a small support set, rather than directly encoding task-specific conditions. The interpolation is guided by patch-wise similarity scores between the query and support conditions, modeled by a matching module [23]. Since image features are inherently task-agnostic, this interpolation-based approach naturally provides a unified representation, enabling effective adaptation across diverse spatial tasks. Furthermore, to facilitate rapid and data-efficient updates, UFC combines episodic meta-learning on a multi-task dataset with a small set of task-specific parameters, enabling effective adaptation under few-shot settings.
Extensive experiments across diverse spatial control modalities, including edge maps, depth maps, normal maps, human pose, and segmentation masks for human body, demonstrate that UFC outperforms prior approaches in few-shot settings by large margins. Notably, UFC achieves fine-grained spatial control using only 30 shots, and even on par with some fully supervised baselines on Normal, Depth, and Canny with just 150 support examples (0.1% of the training data). Furthermore, UFC is compatible with both UNet [42] and DiT [36] backbones, ensuring applicability across recent T2I diffusion architectures. To our knowledge, we propose the first method for few-shot spatial control in text-to-image diffusion models, which enables spatial control image generation with novel conditions at test time using minimal annotated data.
@@ 2 Related Works
Text-to-Image diffusion models [11, 41, 38, 43, 8, 7, 25] have gained widespread adoption for their ability to generate high-quality, semantically aligned images from text prompts. To enable finer control over image structure, various methods [35, 58, 28, 27, 39, 59, 53, 48, 52] proposed incorporating additional spatial control inputs such as edge maps, segmentation masks, or human poses, etc., into diffusion models. ControlNet [58] augments a pretrained T2I diffusion model with a control adapter, initialized from the model itself, to encode spatial conditions and inject them into the frozen diffusion model. Uni-ControlNet [59] extends this by adding lightweight feature extractors to support multiple condition types. Prompt Diffusion [53], built on ControlNet, introduces a vision-language prompting mechanism to enable in-context learning for handling a diverse set of tasks. However, these prior methods are limited to tasks seen during training. To address this limitation, training-free methods [2, 57, 34, 29] have been proposed; but, they often rely on a long process of latent optimization, which substantially increases generation time while providing only limited controllability. Consequently, few-shot adaptation with more fine-grain controllability without incurring significant generation-time overhead remains unexplored.
@@
To extend diffusion models to unseen tasks, several works [5, 60, 55, 46, 12, 26, 20] have explored few-shot learning for image generation. For example, D2C [46] trains a variational autoencoder [24] on a few labeled examples to produce latent representations that condition a diffusion model to generate images in the target domain. Similarly, FSDM [12] uses visual features extracted from a small support set to guide image generation toward unseen classes. While these approaches demonstrate promising few-shot generation capabilities, they primarily focus on domain or class transfer, and are not designed to generalize across diverse spatial control tasks.
@@
Several recent works [13, 56] leverage the analogy between a query source and a support source-target pair to enable generalization in diffusion models on diverse tasks with visual instruction. Analogist [13] treats visual instruction as an inpainting task, arranging query-support pairs in a $2\times 2$ grid and using a diffusion model to complete the missing cell. This training-free method generalizes well to tasks like colorization, deblurring, and editing. While we also leverage query-support analogy in few-shot settings, our goal differs: prior work preserves the support image’s appearance, whereas we aim to generate diverse content guided by spatial conditions.
@@ 3 Background
A text-to-image diffusion model generates an image $x$ conditioned on a text prompt $c_{\text{text}}$ by iteratively denoising a sequence of latent variables. Specifically, the model learns a finite-length Markov chain $\{z_{T-t}\}_{t=0}^{T}$ that starts from standard Gaussian noise $z_{T}\sim\mathcal{N}(0,I)$, where each transition $z_{t+1}\to z_{t}$ progressively removes noise. The final clean variable $z_{0}$ can be either the image itself or a latent representation $z=E(x)$, which is decoded back to the image by $x=D(z)$, where $(E,D)$ denotes an auto-encoder [41, 38]. The chain is the time-reversal of a diffusion process $\{z_{t}\}_{t=0}^{T}$ that gradually corrupts the data by adding small Gaussian noise until it becomes pure noise at $z_{T}$ [47]. To learn each denoising transition, the diffusion network $\mathcal{E}_{\phi}$ is trained to predict the added noise $\epsilon$ at timestep $t$:
@@
where ($z_{0},c_{\text{text}}$) is drawn from the training dataset $\mathcal{D}_{\text{train}}$ and $t$ is uniformly sampled from $\{1,\cdots,T\}$. $z_{t}$ can be computed in closed form from $z_{0}$ and $t$ using the properties of Gaussian variables [16]. After training, an image is generated by running the learned Markov chain from $z_{T}$ down to $z_{0}$.
Although text-to-image diffusion models control the overall content of an image, a text prompt alone is limited in steering the fine-grained spatial structures. Spatial control approaches [35, 58, 28, 48] address this limitation by introducing an additional spatial condition $y_{\tau}$, typically a dense map describing the target image (e.g., edges, depth, or semantic masks). For each condition type (or task) $\tau$, an auxiliary control adapter $\mathcal{G}_{\theta_{\tau}}(y_{\tau})$ [^1] The control adapter can take auxiliary inputs such as noisy latent, timestep, and text prompt, but we omit them for clarity. is trained and injected into a frozen, pre-trained diffusion backbone $\mathcal{E}_{\phi}$, while keeping the backbone parameters $\phi$ fixed to preserve their learned priors:
@@
However, learning a dedicated adapter for each task $\tau$ usually requires thousands of labeled data to achieve reasonable performance [58]. This can be costly if we want to condition the pre-trained model on a novel task $\tau_{\text{novel}}$. On the other hand, reusing an adapter trained on a fixed task(s) for novel tasks often suffers from limited adaptation performance, especially when the new tasks are substantially different from training [59]. Therefore, designing a versatile control adapter that can accommodate arbitrary unseen spatial conditions with a small amount of data remains a key challenge for data-efficient steering of text-to-image diffusion models.
@@ 4.1 Problem Setup
Our goal is to build a universal few‑shot framework that steers a frozen text‑to‑image diffusion model with arbitrary novel spatial conditions, given only a handful of annotated examples (typically a few dozen). To this end, we introduce a universal control adapter $\mathcal{I}(y_{\tau};\mathcal{S}_{\tau})$ that can inject any type of spatial conditions into the frozen diffusion model by directly incorporating the given data as follows:
@@
where $\mathcal{S}_{\tau}=\{(x^{i},y_{\tau}^{i})\}_{i\leq N}$ denotes a support set comprising $N$ image-condition pairs for task $\tau$. We focus exclusively on the design of the adapter $\mathcal{I}$ while the remaining components, such as loss function, diffusion architecture, and injection strategy, are indepedent to the adapter and can be freely chosen from existing spatial control approaches [35, 58, 28].
@@
    Heterogeneous input formats. Spatial conditions vary widely—from extremely sparse cues such as human‑pose keypoints to dense maps such as depth or segmentation. The adapter must transform these disparate inputs into the unified control features that the frozen diffusion model can consume reliably.
@@
    Severe data scarcity. For a new task, we assume only a few dozen labeled examples. To handle distribution shifts without over‑fitting the tiny support set, the adapter needs an adaptation mechanism that is both flexible and efficient.
@@ 4.2 Universal Few-Shot Control Adapter
The patch-wise matching formulation in Eq. 4, inspired by few-shot dense prediction methods [23, 22], offers several advantages for the universal control problem. First, since the task-specific conditions are only used to determine the weights for the task-agnostic visual features, the resulting control features $\mathcal{I}(y_{\tau}^{q};\mathcal{S}_{\tau})$ reside in a unified visual feature space regardless of the task $\tau$. This ensures consistent guidance of the diffusion model and robust generalization to unseen tasks. Second, by composing the control features patch-wise, it can exploit the locality inductive bias on the image-condition pairs and amplify the effective number of support patches (i.e., bases) for each query patch. This provides sufficient representational power to model complex spatial controls from just a few support images. Finally, the general matching rule can wrap any control adapter $G_{\theta_{\tau}}$ by treating it as a condition encoder $g_{\tau}$ and adding an image encoder $f$. UFC can therefore inherit advances in adapter architectures and injection mechanisms developed for spatial control.
To adapt rapidly to a new task without over‑fitting, UFC introduces only a small set of task-specific parameters $\theta_{\tau}$ while sharing the remainder. Starting from a pre-trained visual backbone $g(\cdot;\theta)$, we obtain a condition encoder
@@
where $\theta$ is shared across tasks and $\theta_{\tau}$ is lightweight (e.g., bias [3] or LoRA [18] parameters). After meta-training on a fixed task set (Section 4.4), $\theta$ is frozen and only $\theta_{\tau}$ is fine-tuned for each novel task $\tau$. The image encoder $f$, also initialized from a pre-trained visual backbone, remains frozen throughout to preserve task-agnostic visual features. Leveraging strong priors of large-scale visual pre-training and modern parameter-efficient tuning techniques, UFC generalizes to unseen tasks with minimal task-specific parameters.
@@ 4.3 Architecture
The formulation of UFC’s control adapter in Eq. (4) is general and can be incorporated into various diffusion backbones and adapter architectures. In this section, we describe the instantiation of UFC within the ControlNet framework [58] due to its popularity and strong performance. For clarity, we consider the pre-trained Stable Diffusion [41, 11] with UNet backbone [41] in this section, while the one with DiT backbone [11] is described in Appendix D.
We initialize the condition encoder $g_{\tau}$ and image encoder $f$ as separate copies of the UNet encoder of the pre-trained diffusion model, where the image encoder is kept frozen. For the task-specific parameters $\theta_{\tau}$ of the condition encoder, we adopt bias-tuning [3], which has been proven to be flexible for few-shot learning in dense visual prediction [23, 22]. The matching module is implemented as a multi-head cross-attention block [50], where the timestep embedding is fused through adaptive layer normalization [37, 36] inside the attention mechanism. More implementation details about the matching module can be found in Appendix A.1.
@@ 4.4 Training and Inference
To equip UFC with few-shot learning capabilities, we adopt the standard episodic meta-training protocol, wherein each training iteration is an episode that mirrors the test-time procedure. In each episode, we first sample a specific control task $\tau$ from a meta-training dataset $\mathcal{D}{{}_{\text{train}}}$ that consists of multiple tasks. The sampled task $\tau$ supplies a support set $\mathcal{S}_{\tau}$ and a query set $Q_{\tau}$, where the model exploits the support set to control the diffusion model with the query condition. The model is optimized using the standard denoising loss [16] (or its flow-matching variant [31]):
@@ 5.1 Experimental setup
For experiments using the UNet architecture, we adopt Stable Diffusion v1.5 [41] as the diffusion backbone. Spatial conditions and support images are first projected into the latent space using the pre-trained VAE [24] from the diffusion model, then processed by respective encoders. When encoding these inputs, we fix the denoising timestep to zero, provide empty text prompts, and omit the noisy latent input in order to focus solely on spatial information. Detailed model setting using DiT backbone is provided in the Appendix D.
@@
To enable episodic meta-training, we sample 300K text-image pairs from LAION-400M [45], with 150K containing humans (for Pose and Densepose) and 150K randomly sampled. Images are resized to the resolution $512\times 512$. Spatial conditions are extracted using pretrained off-the-shelf models. We consider six condition modalities that cover diverse semantics: Canny edge [4], HED edge [54], MiDaS Depth and surface Normal [40], human Pose [6], and human Densepose [14]. We consider Densepose as a control signal for semantic segmentation to constrain all segmentation classes to be included in the support set. During the training process, we maintain the setting that the models will be trained on 150K image-condition pairs for each task.
@@
We train UFC (UNet diffusion backbone) for 12.5K iterations with a batch size of 96 on 8 NVIDIA RTX 3090 GPUs, using AdamW [33] with a learning rate of $1\times 10^{-5}$. During inference with the UNet backbone, we adopt the PNDM sampler [32] with 50 denoising steps, classifier-free guidance (CFG) [17] scale of 7.5, and seed 42. For testing, we fine-tune UFC using all image–condition pairs in the support set, but use only a subset of them as conditions during inference to accommodate memory constraints and reduce generation-time overhead.
@@
  • Fully-supervised baselines ControlNet [58] and Uni-ControlNet [59] are both trained in a fully-supervised manner on the entire dataset (150K images for each task). ControlNet is trained separately for each condition type, while Uni-ControlNet is jointly trained with all conditions.
  • Few-shot baselines As the first to tackle spatial control in diffusion models under a few-shot setting, we adapt existing methods, Uni-ControlNet and Prompt Diffusion [53], to align with our few-shot adaptation. For Uni-ControlNet, as mentioned in the original paper, we extend the input layer with additional channels for novel conditions and few-shot fine-tune this layer. For Prompt Diffusion, we evaluate both the zero-shot setup as proposed in the original paper and the full fine-tuning setup. For fair comparison, these few-shot methods are pre-trained using the same meta-training dataset as our method and fine-tuned on the same few-shot data.
  • Training-free baselines We also compare our method with training-free methods, Ctrl-X [29] and FreeControl [34], which directly employ the frozen T2I diffusion models for spatial control. We consider Ctrl-X as our main evaluation but leave comparisons with FreeControl in Appendix B using a smaller evaluation set due to its slow inference speed.
@@
We quantitatively evaluate all methods on COCO 2017 [30] validation split, which contains 5,000 images with associated captions. All images and their corresponding conditions are resized to $512\times 512$. For tasks Pose and Densepose, we restrict evaluation to images containing humans. To assess image generation quality, we report the Fréchet Inception Distance (FID) [15]. To evaluate controllability, we extract spatial conditions from the generated images using the same pre-trained networks used for dataset construction and compare them with the query conditions using task-specific metrics. Specifically, we use the Structural Similarity Index (SSIM) for Canny and HED, Average Precision ($\text{AP}^{50}$) computed using Object Keypoint Similarity for human Pose, Mean Squared Error (MSE) for Depth prediction, Mean Angular Error (MAE) for surface Normal, and mean Intersection over Union (mIoU) for human segmentation in Densepose.